OpenStudy (anonymous):

Negative one times the sum of twice a number and 3 is equal equal to two times the difference of-4 times the number and 3. What is the number?

OpenStudy (anonymous):

write and equation and solve

OpenStudy (anonymous):

hi. are u still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. so what do you think the equation might look like?

OpenStudy (anonymous):

-1*3n=2*(-4*3n)

OpenStudy (anonymous):

i think you are close, but we need you to get closer.

OpenStudy (anonymous):

is it -1(3+n)=2(-4*3n)

OpenStudy (anonymous):

so lets take the left hand side of the equation the first part. it says -1 times the sum of 2 times a number and 3. you are closer. just the left side first.

OpenStudy (anonymous):

equals -3+-n

OpenStudy (anonymous):

i think it would be -1(2n+3)= for the left hand side. that gives -1 times the entire sum of 2n (2 times a number) and 3. the and implies addition

OpenStudy (anonymous):

do you see that?

OpenStudy (anonymous):

i am sorry. the sum implies addition

OpenStudy (anonymous):

please feel free to ask questions

OpenStudy (anonymous):

how did u get the 2n to the left side

OpenStudy (anonymous):

Reread the first part of the problem

OpenStudy (anonymous):

that isnt part of the question

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

wht will be on the right side

OpenStudy (anonymous):

it is ok i know these can be confusing

OpenStudy (anonymous):

wht grade r u in

OpenStudy (anonymous):

sorry, my computer froze

OpenStudy (anonymous):

i am a math teacher

OpenStudy (anonymous):

now, what do you think the right hand side of the equation looks like?

OpenStudy (anonymous):

u still there candence?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes my computer froze 2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so before it happens again, what do you think the right hand side of the equation looks like? give it your best guess

OpenStudy (anonymous):

i thing it is 2-4(3n)

OpenStudy (anonymous):

your clue here is that it is two times the difference. that means that the 2 multiplies the difference. the difference is in the parentheses.

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

so if you can figure out how to write the difference of -4 times n and 3, you can solve for the answer

OpenStudy (anonymous):

2(-4*3-n)

OpenStudy (anonymous):

closer. 2(-4n-3) is the right side

OpenStudy (anonymous):

now put the left and right hand sides together to solve for n.

OpenStudy (anonymous):

that means just put an equals sign in between them

OpenStudy (anonymous):

so it should look like this. -1(2n+3)=2(-4n-3)

OpenStudy (anonymous):

is it 0.5

OpenStudy (anonymous):

yes. that is what i got

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

you still there candence?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it should be -0.5

OpenStudy (anonymous):

your final steps should be -2x-3=-8x-6 and then 6x=-3 and then dividing by 6 gives x=-.5

OpenStudy (anonymous):

yea but i did it different

OpenStudy (anonymous):

-2n-3=-8n-6 +2n +2n -3=-6n-6 +6 +6 3=-6n - --- -6 -6 n=-0.5

OpenStudy (anonymous):

perfect. you got the negative sign in there. good job of showing all your work. if you do problems like that on your tests, your teacher will love you.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

is there anything else i can help you with?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

admission for a water park is $17.50 per day. a season pass cost $125. a locker rental costs $3.50 per day. 1)hat is an equation that represents the relationship the cost of a daily pass and season pass cost? 2)how many days would you have to go to the water park for the season pass to save you money?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any ideas for the cost if you do a daily pass?

OpenStudy (anonymous):

i think the equation is d(17.50+3.50)= 125+3.50d

OpenStudy (anonymous):

you are absolutely correct, yayyy!! remember that the left hand side of your equation is the cost for a daily pass and the right hand side of your equation is the cost of a season pass.

OpenStudy (anonymous):

so then i solve

OpenStudy (anonymous):

for d

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you have an answer for d?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

take your time

OpenStudy (anonymous):

i got 7.1

OpenStudy (anonymous):

so for the first part of the question, the cost using a daily pass is c=d(17.50+3.50) and the cost using a seasonal pass is c=125+3.50d. Be sure you write that too because that answers the first part of the question. right now you are answering the second part.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well, to solve for d, you have what you wrote above, which after using the distributive property and simplifying becomes, 125=17.50d

OpenStudy (anonymous):

i got 7.1

OpenStudy (anonymous):

you may have made a small error. i got 125/7, which means that you would have to use the seasonal pass at least 18 days before it paid off

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

do you see where you turned left instead of right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you got 17. something close to 18

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont have my calculator lol

OpenStudy (anonymous):

nice candence. very nice you did really well.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

i hope to help you again some time. Unfortunately, i must leave now. i hope you have a wonderful evening . take care

OpenStudy (anonymous):

you 2

OpenStudy (anonymous):

:))